OpenStudy (binary3i):

what do you say about the integral of sin(1/x)?

OpenStudy (anonymous):

Write it as a summation: \[\int\limits \left[ \sum_{k=0}^{\infty}\frac{ (-1)^{k} \left( \frac{1}{x} \right) ^{2k+1}}{{(2k+1)!}} \right]dx\]

OpenStudy (anonymous):

Use the alternating series estimatino theorem to estimate the value.

OpenStudy (anonymous):

integral sin(1/x) dx = x sin(1/x)-cos(1/x)+constant Alternate Form: -Ci(1/x)+1/2 i e^(-i/x) x-1/2 i e^(i/x) x+constant Series expansion of the integral at x=0: -2 i pi floor((arg(x)+pi)/(2 pi))+i pi floor((arg(x))/pi+1/2)+O(x^9)+(x+O(x^7)) sin(1/x)+(-x+2 x^3-24 x^5+O(x^7)) sin(1/x)+(x^2-6 x^4+120 x^6+O(x^7)) cos(1/x) Series expansion of the integral at x=∞: (-log(1/x)-gamma+1)+1/(12 x^2)-1/(480 x^4)+1/(30240 x^6)-1/(2903040 x^8)+1/(399168000 x^10)+O((1/x)^11) I hope this answers your question!

OpenStudy (anonymous):

Its not -cos(1/x) its minus the cosine integral. Which is an unsolvable integral. Which must be approximated.

OpenStudy (anonymous):

It is not possible to find the exact value you can approximate the value by putting it inbetween two known integrals....

OpenStudy (anonymous):

By using the alternating series estimation theorem ________.________